Given the quadratic equation: x^2 - 5x + 1 = 0.
A. What is the sum of the roots?

B. What is the product of the roots?

Answer by fractalier(6550)

For x^2 - 5x + 1 = 0,
the sum of the roots is -b/a = 5/1 = 5
and the product of the roots is c/a = 1/1 = 1